OpenStudy (anonymous):

given that (8u-v) perpendicular (4u+3v) and [v]=2[u] find the angle between u and v.

OpenStudy (anonymous):

if two vectors are perpendicular, it means their dot product is zero

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(8u-v)\cdot(4u+3v)=0\\ 32u\cdot u-4u\cdot v+24u\cdot v-3v\cdot v=0\\ 20u\cdot v =3(2u)^2-32u^2\\ 20u\cdot v+20u^2=0\\ 20u\cdot (v+u)=0\\ u=0;\text{OR}u+v=0\\ u=-v \] angle between them is 180°?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

how do u know its 180?

OpenStudy (anonymous):

-ve sign

OpenStudy (anonymous):

and isn't it supposed to be 20u⋅v+26u2=0? not 20

OpenStudy (anonymous):

3*2^2-32=-20

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

so the angle can also be 0 right?

OpenStudy (anonymous):

no. if 0, then sign is +ve

OpenStudy (anonymous):

but how do u know that -ive sign means 180?